When mean optical power launched into an 8 km of fiber is 120 mW, the output mean optical power is 3 microwatt. Determine:
(a) overall loss in dB
(b) single attenuation per kilometre
1
Expert's answer
2012-09-18T11:04:17-0400
a) losses=8*120*10^(-3)/3*10^(6)=32dB
b) 3/8*10^(6)=37.5dB/km